OpenStudy (anonymous):

Can someone help me solve: 3(x-2)^2/4=24

OpenStudy (anonymous):

is it \[\frac{3(x-2)^2}{4}=24\]?

OpenStudy (anonymous):

if so, start with \[(x-2)^2=24\times \frac{4}{3}\] \[(x-2)^2=32\] and then it should be straighforward take the square root, don't forget the \(\pm\) then add two